The height of a cylindrical tank, containing 15 lb of gas at 70 psia and 115o F is triple its diameter; the gas constant R = 45 ft-lb/lboR. Find the tank height. 


1
Expert's answer
2021-12-09T09:15:18-0500

"V=\frac{mRT}p,"

"V=Sh=\frac{\pi d^2}4 h=\frac{\pi h^3}{36},"

"h=\sqrt[3]{\frac{36 mRT}{\pi p}}=7.61~ft."